OpenStudy (anonymous):

how to find (a) the complement and (b) the supplement of an angle at the given measure. 1.30° 2.54° 3.14°21' 4.29°50' 5.20°10'30"

OpenStudy (anonymous):

Tow angles are complementary if their sum is 90 degrees. Two angles are supplementary if their sum is 180 degrees.

OpenStudy (anonymous):

No trigonometry needed. The complement of a 40.6 degree angle is 90-40.6 = 49.4 degrees.

OpenStudy (anonymous):

uhh ok ..

OpenStudy (anonymous):

Same procedure for the supplement of an angle.

OpenStudy (anonymous):

30° x+30°=90 x=30°-90° x=60° complementary

OpenStudy (anonymous):

like this??

OpenStudy (anonymous):

Yes, or just subtract the angle from 90 (or from 180 in the case of the supplement of the angle).

OpenStudy (anonymous):

That is, the complement of any angle, x, is just 90 - x. The supplement of any angle, y, is just 180-y.

OpenStudy (anonymous):

30° x+30°=180 x=30°-180° x=150° supplementary

OpenStudy (anonymous):

x = 180 - 30 (not 30 - 180) (Error in your third line)

OpenStudy (anonymous):

how about 14° 21' ?

OpenStudy (anonymous):

how to solve number 3 ??